# Geometric Approach to Light Rings in Axially Symmetric Spacetimes

### v2

@@ Appendix B Geodesic Curvature in Randers-Finsler Optical Geometry
Then we are able to apply the aforementioned Finslerian geodesic curvature $\kappa_{g}^{(F)}$ to the optical geometry of axially symmetric rotational spacetimes, which gives a Randers-Finsler geometry in expression (7). Since the 2-dimensional Randers-Finsler geometry can be decomposed into a Riemannian geometry part $\alpha$ and a non-Riemannian part $\beta$, it reasonable to anticipate that the Finslerian geodesic curvature $\kappa_{g}^{(F)}$ for a continuous curve is composed of two parts: the geodesic curvature with respect to the Riemannian metric $\alpha_{ij}$, and the additional contribution due to the existence of the non-Riemannian part $\beta$. Therefore, the Finslerian geodesic curvature is given by
@@
For the axially symmetric rotational spacetime, its optical geometry takes the form

### v3

@@ Appendix B Geodesic Curvature in Randers-Finsler Optical Geometry
Then we are able to apply the aforementioned Finslerian geodesic curvature $\kappa_{g}^{(F)}$ to the optical geometry of axially symmetric spacetimes, which gives a Randers-Finsler geometry in expression (7). Since the 2-dimensional Randers-Finsler geometry can be decomposed into a Riemannian geometry part $\alpha$ and a non-Riemannian part $\beta$, it reasonable to anticipate that the Finslerian geodesic curvature $\kappa_{g}^{(F)}$ for a continuous curve is composed of two parts: the geodesic curvature with respect to the Riemannian metric $\alpha_{ij}$, and the additional contribution due to the existence of the non-Riemannian part $\beta$. Therefore, the Finslerian geodesic curvature is given by
@@
For the axially symmetric spacetime, its optical geometry takes the form